The tallest living man at one time had a 248 Why time had a of Heights of men at that time had a man of 175 88 cm and a standard deviation of 759 Which of these two men had the that was more extreme?


1
Expert's answer
2021-03-05T01:12:34-0500

Total problem: The tallest living man at one time had a height of 248 cm. The shortest living man at that time had a height of 65.5 cm. Heights of men at that time had a mean of 175.88 cm and a standard deviation of 7.59 cm. Which of these two men had the height that was more​ extreme?

Z score for the tallest person:

"\frac{248-175.88}{7.59} = 9.50"

Z score for the tallest person:

"\frac{65.5 - 175.88}{7.59} = -14.54"

Since the z score for the tallest man is z=9.50 and the z score for the shortest man is z=-14.54, the shortest man had the more extreme height. Since |14.54|>|9.50|.